According to the table at right, if the fourth and fifth largest firms in the industry merge, the four-firm concentration ratio in the industry will be Firm Annual Sales $1,000 900 O A. 69.0 percent. OB. 82.5 percent. OC. 35.8 percent. OD, 52.5 percent. 120 75 40 800 1,200 1,050 75 600
Answer #1

Here fourth largest firm is B and the fifth largest firm is G. So annual sale after merging of these two firms is (900+800)=$1700

Top four firm's sale is (1700+1200+1050+1000)=$4950

Total sale of (1000+900+120+75+50+40+800+1200+1050+90+75+600)=$6000

Four firm concentration ratio is (4950/6000)*100=82.5%

Option B is correct.